Find the volume generated by revolving the given region about the given axis.

(a) The region bounded by y = x4 , x = 1 and y = 0 about Y axis.

(b) The triangle with vertices (1, 1),(1, 2)(2, 2) about X axis.

(c) The region in the first quadrant bounded by x = y − y3 , x = 1 and y = 1 about X axis. 


1
Expert's answer
2021-11-08T19:50:26-0500

(a)


"A_1(y)=\pi x^2=\pi\sqrt{y}"

"A_2(y)=\pi(x)^2=\pi"

"V=\displaystyle\int_{0}^{1}(\pi-\pi\sqrt{y})dy=\pi[y-\dfrac{2}{3}y^{3/2}]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=\dfrac{\pi}{3} (cubic\ units)"

(b)


"V=\displaystyle\int_{1}^{2}(\pi(2)^2-\pi x^2)dy=\pi[4x-\dfrac{1}{3}x^{3}]\begin{matrix}
 2\\
1
\end{matrix}"

"=\pi(8-\dfrac{8}{3}-(4-\dfrac{1}{3})=\dfrac{5}{3}(cubic\ units)"

(c)



"V=\displaystyle\int_{0}^{1}2\pi y(1-(y − y^3))dy=2\pi[\dfrac{y^2}{2}-\dfrac{y^3}{3}+\dfrac{y^4}{4}]\begin{matrix}
1\\
0
\end{matrix}"

"=\dfrac{5\pi}{12}(cubic\ units)"